# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 3.1 Observables and derivations
    If $U$ and $V$ are transverse, then this is jointly continuous bilinear map from ${\mathfrak{D}}_{al}(U)\times{\mathfrak{D}}_{al}(V)$ to ${\mathfrak{D}}_{al}(U\cap V)$.
In particular since $\{0\}$ and ${\mathbb{R}}^{n}$ are transverse, ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ acts continuously on ${\mathfrak{D}}_{al}(\{0\})$ and this action is easily seen to extend to a continuous action of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on the space ${\mathscr{A}}_{a\ell}$. We denote the action of ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ by ${\mathcal{A}}\mapsto{\mathsf{F}}({\mathcal{A}})$. By Prop. 3.6, it is given by
${\mathsf{F}}({\mathcal{A}})=\sum_{Y\in{\mathbb{B}}_{n}}[{\mathsf{F}}^{Y},{\mathcal{A}}].$ (19)
It is not hard to check that for any $X\in{\mathbb{B}}_{n}$ and any ${\mathcal{A}}\in{\mathscr{A}}(X)$ we have ${\overline{\rm tr}}_{X^{c}}\left({\mathsf{F}}({\mathcal{A}})\right)=[{\mathsf{F}}_{X},{\mathcal{A}}]$ and so the action of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ on ${\mathscr{A}}_{a\ell}$ is faithful. Thus, elements of ${\mathfrak{D}}_{al}(U)$ for any $U\subset{\mathbb{R}}^{n}$ may be identified with a subset of the Fréchet-continuous derivations of ${\mathscr{A}}_{a\ell}$. By Proposition 3.2, the Fréchet-Lie algebra ${\mathfrak{D}}_{al}(\{0\})$ is identified with the Fréchet-Lie algebra ${{\mathfrak{d}}_{al}}$ of traceless anti-hermitian elements of ${\mathscr{A}}_{a\ell}$ acting by inner derivations.
@@ 3.2 Automorphisms
$\frac{d}{dt}{\mathcal{A}}(t)={\mathsf{F}}(t)({\mathcal{A}}(t))$
with the initial condition ${\mathcal{A}}(0)={\mathcal{A}}$ has a unique solution ${\mathcal{A}}(t)\in{\mathscr{A}}_{a\ell}$ for all $t\in{\mathbb{R}}$. Denote by $\alpha^{\mathsf{F}}_{t}$ the map taking ${\mathcal{A}}$ to ${\mathcal{A}}(t)$. It is a continuous automorphism of the Lie algebra ${\mathscr{A}}_{a\ell}$ that preserves the *-operation.
Definition 3.6.
We call any automorphism of the form $\alpha^{\mathsf{F}}_{t}$ for some smooth map ${\mathsf{F}}:{\mathbb{R}}\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ a locally-generated automorphism, or LGA for short.
It is shown in [11] that every LGA extends to a continuous *-automorphism of the quasilocal algebra ${\mathscr{A}}$, and that the set of LGAs forms a group under composition.
@@ 3.3 States
By a state $\psi$ we will mean a state on the quasilocal algebra ${\mathscr{A}}$. If ${\mathsf{F}}$ is an inner derivation (see Def. 3.4), we define its $\psi$-average as the evaluation of $\psi$ on the corresponding element of ${{\mathfrak{d}}_{al}}\subset{\mathscr{A}}_{a\ell}$. The group of LGAs acts on states by pre-composition, which we denote $\psi^{\alpha}:=\psi\circ\alpha$. We say an LGA $\alpha$ preserves a state $\psi$ if $\psi^{\alpha}=\psi$. We say an element ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ preserves $\psi$ if $\psi({\mathsf{F}}({\mathcal{A}}))=0$ for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$, which is equivalent to the one-parameter group of automorphisms $t\mapsto\alpha_{t}^{\mathsf{F}}$ corresponding to a constant map $t\mapsto{\mathsf{F}}$ preserving $\psi$.
Definition 3.7.
For any $U\subset{\mathbb{R}}^{n}$ define ${\mathfrak{D}}_{al}^{\psi}(U)$ as the set of all elements of ${\mathfrak{D}}_{al}(U)$ that preserve $\psi$.
It is easy to check that ${\mathfrak{D}}_{al}^{\psi}(U)$ is a closed subset of ${\mathfrak{D}}_{al}(U)$, and that if ${\mathsf{F}}$ and ${\mathsf{G}}$ preserve $\psi$, then $[{\mathsf{F}},{\mathsf{G}}]$ preserves $\psi$. Thus the analog of Propositions 3.4 and 3.6 and Corollary 3.1 hold for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$. Proposition 3.5 on the other hand, does not hold for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$ for a general state $\psi$. To circumvent this, we will restrict to gapped states, where quasiadiabatic evolution [18, 19, 20] can be used to prove the analog of Proposition 3.5.
Definition 3.8.
A state $\psi$ is gapped if there exists ${\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ and $\Delta>0$ such that for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ one has
@@
Remark 3.1.
The meaning of this condition becomes more transparent if one recalls that any ${\mathsf{H}}\in{\mathfrak{D}}_{al}$ is a generator of a one-parameter group of *-automorphisms of ${\mathscr{A}}$ [11]. The condition (20) implies that $\psi$ is invariant under this one-parameter group of automorphisms [21], and that the corresponding one-parameter group of unitaries in the GNS representation of ${\mathscr{A}}$ has a generator whose spectrum in the orthogonal complement to the GNS vacuum vector is contained in $[\Delta,+\infty)$. The condition (20) also implies that $\psi$ is pure [22].
Remark 3.2.
If $\psi$ is a gapped state of ${\mathscr{A}}$ and $\alpha$ is an LGA, then $\psi^{\alpha}$ is also a gapped state. In [11] it was proposed to define a gapped phase as an orbit of gapped state under the action of the group of LGAs.
@@
Proposition 3.7.
Suppose $\psi$ is gapped, the corresponding Hamiltonian is ${\mathsf{H}}$. Then there are linear functions
$\mathcal{J}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}^{\psi}({\mathbb{R}}^{n})$
$\mathcal{K}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$
@@
    For every $k>0$, $U\subset{\mathbb{R}}^{n}$, and ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ we have
    $\|\mathcal{J}({\mathsf{F}})\|_{U,k}$ $\leq C_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$
    $\|\mathcal{K}({\mathsf{F}})\|_{U,k}$ $\leq C^{\prime}_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$
@@
    ${\mathsf{F}}=\mathcal{J}({\mathsf{F}})-\mathcal{K}([{\mathsf{H}},{\mathsf{F}}]).$
Using the above Lemma we will prove the analog of Proposition 3.5 for the spaces ${\mathfrak{D}}_{al}^{\psi}(U)$.
Proposition 3.8.
For any $U,V\subset{\mathbb{R}}^{n}$ consider the sequence
[…]
where $\alpha({\mathcal{A}})=({\mathcal{A}},-{\mathcal{A}})$ and $\beta({\mathcal{A}},{\mathcal{B}})={\mathcal{A}}+{\mathcal{B}}$.
@@
To prove Proposition 3.8 we will need the following geometric result
Lemma 3.6.
Let $U,V\subset{\mathbb{R}}^{n}$ and define $U^{\prime}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\}$. Then $U^{\prime}$ and $U\cup V$ are transverse and their intersection is $U$.
@@
The proof of Proposition 3.5 goes through unmodified except for the definition of $\gamma$, which needs to be changed to ensure that the image of $\gamma$ consists of derivations that preserve $\psi$. Suppose $U,V\subseteq{\mathbb{R}}^{n}$ and ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U\cup V)$. Define
$U^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\},$
$V^{\prime}$ :=$:=\{x\in{\mathbb{R}}^{n}:d(x,V)\leq d(x,U)\}.$
@@
Let $\gamma^{U,V}$ (resp. $\gamma^{U^{\prime},V^{\prime}}$) be the splitting from Proposition 3.5 with the sets $U$ and $V$ (resp. $U^{\prime}$ and $V^{\prime}$). Define $\tilde{\gamma}=(\tilde{\gamma}_{1},\tilde{\gamma}_{2})$ as
$\tilde{\gamma}_{i}({\mathsf{F}})$ $:=\mathcal{J}(\gamma_{i}^{U,V}({\mathsf{F}}))-\mathcal{K}([\mathcal{J}(\gamma_{i}^{U^{\prime},V^{\prime}}({\mathsf{H}})),{\mathsf{F}}])$

### v3

@@ 2.1 Observables and derivations
    If $U$ and $V$ are transverse, then this is a jointly continuous bilinear map from ${\mathfrak{D}}_{al}(U)\times{\mathfrak{D}}_{al}(V)$ to ${\mathfrak{D}}_{al}(U\cap V)$.
Example 35.
Consider a system of fermionic particles as in Example 19. UAL Hamiltonians of the form (25) form a Lie subalgebra of ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$. When expressed in terms of infinite matrices, the Lie bracket is the matrix commutator: ${\mathsf{F}}(h),{\mathsf{F}}(h^{\prime})]={\mathsf{F}}([h,h^{\prime}])$. The derivation ${\mathsf{F}}(h)$ belongs to ${\mathfrak{D}}_{al}(U)$ if $h(x,y)$ decays rapidly when either $x$ or $y$ are far from $U$. That is,
$\sup_{x,y\in\Lambda}\|h(x,y)\|(1+d(x,y)+d(x,U)+d(y,U))^{\ell}<\infty,\quad\forall\ell\geq 0.$ (69)
Almost-local observables and anchored derivations
Earlier in this section, UAL derivations were defined as unbounded derivations with domain the space $\mathscr{A}$$_{\ell}$ of strictly local observables. This was made necessary by the fact that in general a UAL derivation cannot be continuously extended to the entire space $\mathscr{A}$} of quasilocal observables. It is natural to ask whether the use of unbounded operators can be avoided by finding a suitable subalgebra of $\mathscr{A}$} such that UAL derivations act as bona-fide derivations on this algebra. It is indeed possible [19], and below we define the space ${\mathscr{A}}_{a\ell}$ of almost-local observables, equip it with a Fréchet topology, and show that UAL derivations act as continuous derivations of ${\mathscr{A}}_{a\ell}$.
Pick any $o\in{\mathbb{Z}}^{n}$ and for each $k\geq 0$ define the following norm on $\mathscr{A}$$_{\ell}$:
$\|\mathcal{A}\|_{k}:=\|\mathcal{A}\|+\sup_{r\in{\mathbb{N}}\cup\{0\}}(1+r)^{k}\|\mathcal{A}-{\overline{\rm tr}}_{B_{o}(r)^{c}}(\mathcal{A})\|.$ (70)
An elementary argument (see [19]) shows that the norms resulting from different choices of $o\in{\mathbb{Z}}^{n}$ are equivalent. [^5] In [19] a slightly different set of norms was used: $o$ was taken to be a lattice site, while $r$ was allowed to be an non-negative real number. The norms we use in this paper are equivalent but more convenient since $B_{o}(r)$ is a brick for all allowed choices of $o$ and $r$.
Definition 36.
Let ${\mathscr{A}}_{a\ell}$ be the completion of $\mathscr{A}$$_{\ell}$ in the topology generated by the norms $\|\cdot\|_{k}$ for $k\geq 0$.
This is the space of almost-local observables. It is shown in [19, Appendix B] that it is a subalgebra of $\mathscr{A}$}. It is dense because it contains the algebra $\mathscr{A}$$_{\ell}$ of strictly local operators.
Proposition 37.
Suppose $U\subseteq{\mathbb{R}}^{n}$ is non-empty and bounded and let ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$. Then the sum $\sum_{X\in{\mathbb{B}}_{n}}{\mathsf{F}}^{X}$ is norm-absolutely convergent and defines an embedding
$\Sigma:{\mathfrak{D}}_{al}(U)\hookrightarrow{\mathscr{A}}_{a\ell}$ (71)
whose image is the subspace of traceless even anti-hermitian elements of ${\mathscr{A}}_{a\ell}$.
To prove Proposition 37 we introduce the following Lemma
Lemma 38.
Let $X\in{\mathbb{B}}_{n}$, $o\in{\mathbb{R}}^{n}$, and $r>0$.
  1. i)
    $1+r>1+\operatorname{diam}(X)+d(o,X)\implies X\subseteq B_{o}(r)$
  2. ii)
    $1+r\leq\frac{1}{4}(1+\operatorname{diam}(X)+d(o,X))\implies X\nsubseteq B_{o}(r)$
Proof.
The first statement follows from the inequality $d(o,x)\leq d(o,X)+\operatorname{diam}(X)$ for every $x\in X$. For the second, let $x_{1},x_{2}\in{\mathbb{R}}^{n}$ be two points with $d(x_{1},x_{2})=\operatorname{diam}(X)$. By the triangle inequality, we have $d(o,x_{1})+d(o,x_{2})\geq\operatorname{diam}(X)$ so without loss of generality $d(o,x_{1})\geq\operatorname{diam}(X)/2$. Since $d(o,x_{1})\geq d(o,X)$, we thus have
$d(o,x_{1})$ $\geq\frac{1}{2}\left(\frac{\operatorname{diam}(X)}{2}+d(o,X)\right)$
$>\frac{1}{4}(1+\operatorname{diam}(X)+d(o,X))-1,$ (72)
which proves ii)
We now prove Proposition 37:
Proof of Proposition 37.
We can assume without loss of generality that $U=\{o\}$ for some $o\in{\mathbb{R}}^{n}$. We have
$\sum_{Y\in{\mathbb{B}}_{n}}\|{\mathsf{F}}^{Y}\|$ $\leq\|{\mathsf{F}}\|_{\{o\},2n+2}\sum_{Y\in{\mathbb{B}}_{n}}(1+\operatorname{diam}(Y)+d(Y,o))^{-2n-2},$ (73)
so Lemma 8 the above sum is finite. To show that it is injective, notice that for any $X\in{\mathbb{B}}_{n}$ we have $\pi^{X}(\Sigma({\mathsf{F}}))={\mathsf{F}}^{X}$, and so $\Sigma({\mathsf{F}})=0$ implies ${\mathsf{F}}^{X}=0$ for all $X\in{\mathbb{B}}_{n}$. To show that $\Sigma$ is an embedding, we will show that for any $k\geq 0$ there are constants $C_{k},C_{k}^{\prime}>0$ such that
$\|\Sigma({\mathsf{F}})\|_{k}\leq C_{k}\|{\mathsf{F}}\|_{\{o\},k+2n+2}$ (74)
and
$\|{\mathsf{F}}\|_{\{o\},k}\leq C_{k}^{\prime}\|\Sigma({\mathsf{F}})\|_{k}.$ (75)
From the equality
$\Sigma({\mathsf{F}})-{\overline{\rm tr}}_{B_{o}(r)^{c}}(\Sigma({\mathsf{F}}))=\sum_{X\nsubseteq B_{o}(r)}{\mathsf{F}}^{X},$ (76)
we get for any $r>0$ and any $k\geq 0$:
$(1+r)^{k}\|\Sigma({\mathsf{F}})-{\overline{\rm tr}}_{B_{o}(r)^{c}}(\Sigma({\mathsf{F}})\|$
$\leq$ $(1+r)^{k}\sum_{X\nsubseteq B_{o}(r)}\|{\mathsf{F}}^{X}\|$
$\leq$ $\|{\mathsf{F}}\|_{\{0\},k+2n+2}(1+r)^{k}\sum_{X\nsubseteq B_{o}(r)}(1+\operatorname{diam}(X)+d(o,X))^{-k-2n-2}$
$\leq$ $\|{\mathsf{F}}\|_{\{o\},k+2n+2}\sum_{X\nsubseteq B_{o}(r)}(1+\operatorname{diam}(X)+d(o,X))^{-2n-2}$
$\leq$ $C_{k}\|{\mathsf{F}}\|_{\{o\},k+2n+2}$ (77)
where we used Lemma 38 i) in the second-last line and Lemma 8 in the last line. For the second inequality, using (76) and Lemma 11 we have
$\|{\mathsf{F}}^{X}\|$ $=\|\pi^{X}(\Sigma({\mathsf{F}})-{\overline{\rm tr}}_{B_{o}(r)^{c}}(\Sigma({\mathsf{F}}))\|\leq 4^{n}\|\Sigma({\mathsf{F}})-{\overline{\rm tr}}_{B_{o}(r)^{c}}(\Sigma({\mathsf{F}})\|$ (78)
whenever $X$ is not contained in $B_{o}(r)$. Lemma 38 ii) guarantees this for $1+r=\frac{1}{4}(1+\operatorname{diam}(X)+d(o,X))$. Thus (78) gives
$(1+\operatorname{diam}(X)+d(o,X))^{k}\|{\mathsf{F}}^{X}\|$ $\leq 4^{n}(1+\operatorname{diam}(X)+d(o,X))^{k}\|\Sigma({\mathsf{F}})-{\overline{\rm tr}}_{B_{o}(r)^{c}}(\Sigma({\mathsf{F}})\|$
$=4^{n+k}(1+r)^{k}\|\Sigma({\mathsf{F}})-{\overline{\rm tr}}_{B_{o}(r)^{c}}(\Sigma({\mathsf{F}})\|$
$\leq 4^{n+k}\|\Sigma({\mathsf{F}})\|_{k},$ (79)
which proves (75). Finally, the last statement of the Proposition follows from the decomposition $\mathcal{A}=\sum_{X}\pi^{X}(\mathcal{A})$ for any traceless, even, anti-hermitian element $\mathcal{A}\in{\mathscr{A}}_{a\ell}$. ∎
Proposition 37 shows that ${\mathfrak{D}}_{al}(U)$ for a bounded set $U$ coincides with the set of inner derivations of ${\mathscr{A}}_{a\ell}$:
Definition 39.
For ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$, the following are equivalent:
  1. i)
    ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ for a bounded set $U$
  2. ii)
    There is a $\mathcal{A}\in{\mathscr{A}}_{a\ell}$ such that ${\mathsf{F}}(\mathcal{B})=[\mathcal{A},\mathcal{B}]$ for every $\mathcal{B}\in{\mathscr{A}}_{a\ell}$
A derivation satisfying this condition is called anchored.
@@ 2.2 Automorphisms
$\frac{d}{dt}{\mathcal{A}}(t)={\mathsf{F}}(t)({\mathcal{A}}(t))$ (80)
with the initial condition ${\mathcal{A}}(0)={\mathcal{A}}$ has a unique solution ${\mathcal{A}}(t)\in{\mathscr{A}}_{a\ell}$ for all $t\in{\mathbb{R}}$. While Ref. [19] only considered the case when all $V_{x}$ are even, the proof applies word for word in the ${\mathbb{Z}}_{2}$-graded case. Denote by $\alpha^{\mathsf{F}}_{t}$ the map that takes $\mathcal{A}$} to ${\mathcal{A}}(t)$. It is a one-parameter family of continuous *-automorphisms of the algebra ${\mathscr{A}}_{a\ell}$. By definition, for every ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ it satisfies the differential equation
$\frac{d}{dt}\alpha_{t}^{\mathsf{F}}({\mathcal{A}})=\alpha_{t}^{\mathsf{F}}({\mathsf{F}}(t)({\mathcal{A}})),$ (81)
as well as the initial condition $\alpha_{0}^{\mathsf{F}}={\rm id}$.
Definition 40.
We call any automorphism of the form $\alpha^{\mathsf{F}}_{1}$ for some smooth map ${\mathsf{F}}:[0,1]\rightarrow{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ a locally-generated automorphism, or LGA for short.
It is shown in [19], Section 3.5, that the set of LGAs forms a group under composition.
There is a natural notion of a smooth path between two LGAs $\alpha_{0}$ and $\alpha_{1}$. Namely, a smooth path of LGAs is a 1-parameter family of automorphisms $\alpha_{t}^{\mathsf{F}}$, $t\in[0,1]$ and a smooth path of UAL derivations ${\mathsf{F}}:[0,1]\rightarrow{\mathfrak{D}}_{al}$ which satisfy the equation (81) for all ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}$ and the conditions $\alpha_{0}^{\mathsf{F}}=\alpha_{0}$, $\alpha_{1}^{\mathsf{F}}=\alpha_{1}$. By the results of Section 3.5 of [19], there is a smooth path between any two LGAs.
@@ 2.3 States
By a state $\psi$ of a quantum lattice system we will mean a state of the quasilocal algebra $\mathscr{A}$} which vanishes on ${\mathscr{A}}^{odd}$. If $\mathsf{F}$} is an anchored derivation (see Def. 39), we define $\psi({\mathsf{F}}):=\psi(\Sigma({\mathsf{F}}))$. The group of LGAs acts on states by pre-composition, which we denote $\psi^{\alpha}:=\psi\circ\alpha$. We say an LGA $\alpha$ preserves a state $\psi$ if $\psi^{\alpha}=\psi$. We say an element ${\mathsf{F}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ preserves $\psi$ if $\psi({\mathsf{F}}({\mathcal{A}}))=0$ for any ${\mathcal{A}}\in{\mathscr{A}}_{a\ell}^{even}$, which is equivalent to the one-parameter group of automorphisms $t\mapsto\alpha_{t}^{\mathsf{F}}$ corresponding to a constant map $t\mapsto{\mathsf{F}}$ preserving $\psi$.
Definition 41.
For any $U\subseteq{\mathbb{R}}^{n}$ define ${\mathfrak{D}}^{\psi}_{al}(U)$ as the set of all elements of ${\mathfrak{D}}_{al}(U)$ that preserve $\psi$.
It is easy to check that ${\mathfrak{D}}^{\psi}_{al}(U)$ is a closed subset of ${\mathfrak{D}}_{al}(U)$, and that if $\mathsf{F}$} and $\mathsf{G}$} preserve $\psi$, then ${\mathsf{F}},{\mathsf{G}}]$ preserves $\psi$. Thus the analog of Propositions 28 and 30 and Corollary 34 hold for the spaces ${\mathfrak{D}}^{\psi}_{al}(U)$. Proposition 29, on the other hand, need not hold for the spaces ${\mathfrak{D}}^{\psi}_{al}(U)$ for a general state $\psi$.
The following definition made in [19] is not essential for this paper, but is helpful to keep in mind.
Definition 42.
A smooth path of states is a 1-parameter family of states of the form $\psi_{u}=\omega\circ\alpha(u)$, where $\omega$ is a state and $u\mapsto\alpha(u)$ is a smooth path of LGAs.
In this paper, we will restrict our attention to gapped states, where quasiadiabatic evolution [16, 21, 31] can be used to prove the analog of Proposition 29.
Definition 43.
A state $\psi$ is gapped if there exists ${\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ and $\Delta>0$ such that for any ${\mathcal{A}}\in{\mathscr{A}}_{\ell}$ one has
@@
Remark 44.
The meaning of this condition becomes more transparent if one recalls that any ${\mathsf{H}}\in{\mathfrak{D}}_{al}$ is a generator of a one-parameter group of *-automorphisms of $\mathscr{A}$}. The condition (82) implies that $\psi$ is invariant under this one-parameter group of automorphisms, see [11], Section 5.3.3. Moreover, the corresponding one-parameter group of unitaries in the GNS representation of $\mathscr{A}$} has a generator whose spectrum is contained in $\{0\}\cup[\Delta,+\infty)$. The interval $(0,\Delta)$ is called the spectral gap. The condition (82) also implies that $\psi$ is pure [20].
Remark 45.
In the definition of a smooth family of states, if $\omega$ is gapped, so is $\psi_{u}$ for every $u\in[0,1]$. Indeed, if $\omega$ is gapped with respect to ${\mathsf{H}}\in{\mathfrak{D}}_{al}$, then $\psi_{u}$ is gapped with respect to $\alpha(u)^{-1}({\mathsf{H}})\in{\mathfrak{D}}_{al}$, for the same value of $\Delta>0$.
Remark 46.
In [19] it was proposed to define a gapped phase as an orbit of a gapped state under the action of the group of LGAs. By virtue of Definition 42 and Remark 45, two gapped states of a quantum lattice system are in the same phase iff they can be connected by a smooth path of states.
Remark 47.
The canonical state $\overline{\rm tr}$} is not gapped for any ${\mathsf{H}}\in{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$. Indeed, let ${\mathcal{A}}\in{\mathscr{A}}_{\ell}$. Using the fact that $\overline{\rm tr}$} is a tracial state, one can easily check that the left-hand side of (82) is replaced with its opposite when $\mathcal{A}$} is replaced with $\mathcal{A}$$^{*}$.
Example 48.
Consider a system of fermionic particles (Example 6) with all $H_{x}$ isomorphic to $H$. Let $P:\Lambda\times\Lambda\rightarrow\mathrm{Hom}_{\mathbb{C}}(H,H)$ be a function parameterizing a UAL Hamiltonian as in Example 19. Let ${\mathsf{F}}(P)$ be the corresponding UAL Hamiltonian. Suppose that $P$ is a projector when regarded as an operator on $\ell^{2}(\Lambda,H)$. For any $f\in\ell^{2}(\Lambda,H)$ let $c(f)=\sum_{x,\alpha}f_{\alpha}(x)c_{\alpha,x}$. There is a unique state $\psi$ of the CAR algebra satisfying
$\psi(c(f)^{*}c(f))$ $=0,\quad\forall f\in{\rm im}\,P,$ (83)
$\psi(c(f)c(f)^{*})$ $=0,\quad\forall f\in{\rm im}\,(1-P).$ (84)
This state of the CAR algebra is known as a Fock state, see [11] for details. One can show that $\psi$ is a gapped state with gap $\Delta=1$ for the Hamiltonian ${\mathsf{H}}={\mathsf{F}}(P)$. One can show that a UAL derivation of the form (25) belongs to ${\mathfrak{D}}^{\psi}_{al}$ if and only if the corresponding operator $h$ commutes with $P$.
@@
Proposition 49.
Suppose $\psi$ is gapped, and the corresponding Hamiltonian is $\mathsf{H}$}. Then there are linear functions
$\mathcal{J}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$ (85)
$\mathcal{K}$ $:{\mathfrak{D}}_{al}({\mathbb{R}}^{n})\to{\mathfrak{D}}_{al}({\mathbb{R}}^{n})$ (86)
@@
    For every $k>0$, $U\subseteq{\mathbb{R}}^{n}$, and ${\mathsf{F}}\in{\mathfrak{D}}_{al}(U)$ we have
    $\|\mathcal{J}({\mathsf{F}})\|_{U,k}$ $\leq C_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$ (87)
    $\|\mathcal{K}({\mathsf{F}})\|_{U,k}$ $\leq C^{\prime}_{k}\|{\mathsf{F}}\|_{U,k+4n+3}$ (88)
@@
    ${\mathsf{F}}=\mathcal{J}({\mathsf{F}})-\mathcal{K}([{\mathsf{H}},{\mathsf{F}}]).$ (89)
Using the above Proposition we will prove the analog of Proposition 29 for the spaces ${\mathfrak{D}}^{\psi}_{al}(U)$.
Proposition 50.
For any $U,V\subseteq{\mathbb{R}}^{n}$ consider the sequence
${\mathfrak{D}}^{\psi}_{al}(U\cap V)\xrightarrow{\alpha}{\mathfrak{D}}^{\psi}_{al}(U)\oplus{\mathfrak{D}}^{\psi}_{al}(V)\xrightarrow{\beta}{\mathfrak{D}}^{\psi}_{al}(U\cup V)\to 0,$ (90)
@@
Remark 51.
This Proposition is the only place where we use the existence of the Hamiltonian with respect to which $\psi$ is gapped. In particular, the precise form of this Hamiltonian does not matter, provided it is in ${\mathfrak{D}}_{al}({\mathbb{R}}^{n})$.
To prove Proposition 50 we will need the following geometric result
Lemma 52.
Let $U,V\subseteq{\mathbb{R}}^{n}$ be closed and define $U^{\prime}:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\}$. Then $U^{\prime}$ and $U\cup V$ are transverse and their intersection is $U$.
@@
The proof of Proposition 29 goes through unmodified except for the definition of $\gamma$, which needs to be changed to ensure that the image of $\gamma$ consists of derivations that preserve $\psi$. Suppose $U,V\subseteq{\mathbb{R}}^{n}$ and ${\mathsf{F}}\in{\mathfrak{D}}^{\psi}_{al}(U\cup V)$. Without loss of generality we may assume that $U$ and $V$ are closed. Define
$U^{\prime}$ $:=\{x\in{\mathbb{R}}^{n}:d(x,U)\leq d(x,V)\},$ (94)
$V^{\prime}$ $:=\{x\in{\mathbb{R}}^{n}:d(x,V)\leq d(x,U)\}.$ (95)
@@
Let $\gamma^{U,V}$ (resp. $\gamma^{U^{\prime},V^{\prime}}$) be the splitting from Proposition 29 with the sets $U$ and $V$ (resp. $U^{\prime}$ and $V^{\prime}$). Define $\tilde{\gamma}=(\tilde{\gamma}_{1},\tilde{\gamma}_{2})$ as
$\tilde{\gamma}_{i}({\mathsf{F}})$ $:=\mathcal{J}(\gamma_{i}^{U,V}({\mathsf{F}}))-\mathcal{K}([\mathcal{J}(\gamma_{i}^{U^{\prime},V^{\prime}}({\mathsf{H}})),{\mathsf{F}}])$ (96)